Idiosyncratic Risk

The risk of a security that is associated with random events.

The risk that is specific to one particular investment.

All of these.

None of these.

The risk that can be eliminated by proper diversification.

Alaia stock just paid a dividend of $1.20 and the dividend is expected to remain constant. The stock’s required rate of return is 4%. The current risk-free rate is 3%. What is intrinsic value of the stock today?

P subscript 0 equals fraction numerator $ 1.20 asterisk times open parentheses 1.03 close parentheses over denominator 0.04 minus 0.03 end fraction

P subscript 0 equals fraction numerator $ 1.20 asterisk times open parentheses 1 plus 0.04 close parentheses over denominator 0.04 end fraction

P subscript 0 equals fraction numerator $ 1.20 over denominator 0.04 end fraction

P subscript 0 equals fraction numerator $ 1.20 over denominator 0.04 minus 0.03 end fraction
Answer #1

1. Idiosyncratic risk :

Answer: All of these.

2. Alaia stock just paid a dividend of $1.20 and the dividend is expected to remain constant. The stock’s required rate of return is 4%. The current risk-free rate is 3%. What is intrinsic value of the stock today?

Answer: P subscript 0 equals fraction numerator $ 1.20 over denominator 0.04 end fraction.

Intrinsic value of the stock today = $ 1.20 / 0.04 = $ 30.

As per the dividend discount model,

Stock Price = D 0 ( 1 + g ) / ( r - g ), where g is the growth rate in dividends.

As the dividend is expected to remain constant, g = 0.

Therefore, current stock price = $ 1.20 ( 1 + 0 ) / ( 0.04 - 0 ) = $ 1.20 / 0.04 = $ 30